OpenStudy (anonymous):

Now simplify: 2√17 + 3√7 Im a bit confused, help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

um

OpenStudy (anonymous):

read it

OpenStudy (anonymous):

I dont find anything about square roots...

OpenStudy (anonymous):

help

OpenStudy (anonymous):

o sorry then

OpenStudy (cggurumanjunath):

you can't simplify it further ! it is already simplified.

OpenStudy (anonymous):

the square root of 17. 4 the square root of 17. 2 the square root of 51. 8 the square root of 51.

OpenStudy (anonymous):

choices

OpenStudy (cggurumanjunath):

then there is a mistake in your q ! check the q again!

OpenStudy (cwrw238):

yes its probably 2 terms with sqrt 7 or 2 with sort 17 - also if the options are correct the other numbers are wrong as well

OpenStudy (anonymous):

the square root of 153. - the square root of 17. + the square root of 68.?

OpenStudy (anonymous):

help fast! choices are above

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

A, B, C, or D?

OpenStudy (cwrw238):

hint 153 = 17* 9 68 = 17 * 4

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

D:

OpenStudy (cwrw238):

sqrt 153 = 3 * sqrt17 sqrt 68 = 2 * sqrt17

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

:S

OpenStudy (anonymous):

yea, idk what to do after that. im stck there

OpenStudy (anonymous):

please, i gtg soon. I must solve this

OpenStudy (cwrw238):

3*sqr17 - sqrt17 + 2*sqrt17 = ?

OpenStudy (cwrw238):

just like 3x - x + 2x = 4x

OpenStudy (anonymous):

4 sqrrt17?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

can u do one more quick?

OpenStudy (cwrw238):

hope you've learnt something - study the above

OpenStudy (cwrw238):

i'll point you in right direction

OpenStudy (anonymous):

I did, just forgot the 17 xP

OpenStudy (anonymous):

Can u do one more?

OpenStudy (cwrw238):

as i said i'l help you get there

OpenStudy (anonymous):

its here if u want (04.01) Simplify square root of 8 divided by fifth root of 8 8 to the power of negative 7 over 10 8 to the power of negative 1 over 4 8 to the power of 3 over 10 8 to the power of 7 over 10

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

i gtg in a min ._. this has to be fast ;(

OpenStudy (cwrw238):

b^(½) / 8 ^(1/5) = 8 ^( ½ - 1/5)

OpenStudy (cwrw238):

you subtract the indices so now you need to calculate 1 /2 - 1/5

OpenStudy (cwrw238):

* i meant 8^(½) not b^(½)